OpenStudy (anonymous):

What is the simplified form of (x-2)/(x^2+x-6)÷(x^2+5x+4)/(x+4) 1/(x+3)(x+4) 1/(x+3)(x-2) 1/(x+4)(x-2) 1/(x+3)(x+1)

OpenStudy (anonymous):

I think the answer might be B

OpenStudy (micahm):

Search Results 1. What is the simplified form of (x – 2)(2x +3)? Use the ... https://au.answers.yahoo.com/question/index?qid... Yahoo! Feb 15, 2012 - What is the simplified form of (4p – ...show more ... 2x^2-x-6 12x2 – x – 6 4p2 – 18p + 8 (24x^2-10x-4)*pi --> formula given ... 6x^3+14x^2+5x-3 What is the simplified form of (x-2)/(x2+x-6) divided by (x2+5x+4 ... https://answers.yahoo.com/question/index?qid... Jul 31, 2012 - (x−2)/(x²+x−6) ÷ (x²+5x+4)/(x+4) = (x−2)/(x²+x−6) * (x+4)/(x²+5x+4) = (x−2)/((x−2)(x+3)) * (x+4)/((x+4)(x+1)) = 1/((x+3)(x+1)). 1. What is the simplified form of (x – 2)(2x +3)? Use ... 1 answer Feb 8, 2013 Multiplying and Dividing Rational Expressions .. Two ... 1 answer Apr 10, 2012 What is the simplified form of x+1/x^2+x-6 divide by ... 1 answer Feb 3, 2012 How do i solve x^2-x-6 over x^2-5x+6 in simplest ... 4 answers May 4, 2011 More results from answers.yahoo.com